Circular No. 6835

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1998S AND 1998T Y. L. Qiu, Beijing Astronomical Observatory, reports that the discov-

eries of SN 1998S and SN 1998T were made from CCD frames taken on Mar. 2.68 and Mar. 2.67 UT, respectively (not Mar. 3, as reported on IAUC 6829 and 6830). CCD images taken on Feb. 23.7 show no star at the position of SN 1998S; images taken on Feb. 7.8 show no star at the position of SN 1998T. Magnitude estimates (visual unless otherwise noted) for SN 1998S: Mar. 7.69 UT, 12.6 (S. Yoshida, Waseda University, Japan); 9.91, 12.5 (J. C. Merlin, Le Creusot, France; CCD, using GSC stars); 10.18, 12.8 (P. Schmeer, Bischmisheim, Germany). Corrigendum. On IAUC 6832, line 10, for A.J. 289, 52) read Ap.J. 289, 52)

AX J0049–729 J. Yokogawa and K. Koyama, Kyoto University, report: “An ASCA

observation made on 1997 Nov. 13 revealed a new x-ray pulsar at α = 0^{h}49^{m}01^{s}, δ = −72^{o}51′46′′ (equinox 2000.0; ± 2′ at 90-percent confidence) with a pulsation period of 74.675 ± 0.006 s. This would seem to be the same source having the pulsation period of 74.8 ± 0.4 s in the RXTE 2^{o} field-of-view centered on SMC X-3 (IAUC 6803). The x-ray flux in the band 0.7–10 keV was 1.2 × 10^{−11} erg s^{−1} cm^{−2}, with sinusoidal pulse modulation. The spectrum was flat, with a power law having a photon index of 0.9 ± 0.1.”

IRAS 18325–5926 A. C. Fabian, J. C. Lee, K. Iwasawa, Institute of Astronomy, Univer-

sity of Cambridge; W. N. Brandt, Department of Astronomy and Astrophysics, Pennsylvania State University; and C. S. Reynolds, JILA, University of Colorado, report the presence of periodicities of 11 and 7.8 hr in RXTE observations of the Seyfert galaxy IRAS 18325–5926 on 1997 Dec. 25–27 and 1998 Feb. 21–23, respectively: “These are significantly shorter than the period found in ASCA data from 1997 March of 16 hr

(see http://xxx.lanl.gov/abs/astro-ph/?9801226). The power spec-

trum of the February lightcurve also shows other strong peaks. It is likely that these are due to quasiperiodic oscillations in this massive system, similar to those found in some Galactic blackhole candidates. The changing period is also consistent with the hypothesis of a binary blackhole merger event, which would then occur on or before 1998 Apr. 21. Despite the low a priori probability for such an event, we report it for its potential importance. We urge further observations at all wavelengths.”

1998 March 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green